OpenStudy (anonymous):

What are the x- and y-intercepts of y = 2x^2 - 10x + 8

OpenStudy (anonymous):

y intercept = 8

OpenStudy (anonymous):

let x=0 to find the y-int. what do you get?

OpenStudy (anonymous):

let y=0 to find the x-int. you will have to factorize and you will get two intercepts for x

OpenStudy (anonymous):

0=2x^2-10x+8 0=x^2-5x+4 0=(x-1)(x-4) x=1 and x=4

OpenStudy (anonymous):

can you factor the problem first?

OpenStudy (anonymous):

2 (x^2 -5+4)=y

OpenStudy (anonymous):

make it more do able

OpenStudy (anonymous):

factor it again of course to your liking

OpenStudy (anonymous):

either way is fine :)

OpenStudy (anonymous):

:)